# InfiniteHBD: Building Datacenter-Scale High-Bandwidth Domain for LLM with Optical Circuit Switching Transceivers

### v3

@@ 3. Design Overview
Design 1: Silicon Photonics based OCS transceiver (OCSTrx) (§4.1). To enable large-scale deployment, we require a low-cost, low-power transceiver with Optical Circuit Switching (OCS) support. Unlike prior high-radix switches solutions that rely on MEMS-based switching (Urata et al., 2022a; Yeow et al., 2001), we leverage advances in Silicon Photonics (SiPh), which offer a simpler structure, lower cost, and reduced power consumption—making them well-suited for commercial transceivers.
Our SiPh-based OCS transceiver (OCSTrx), shown on the left of Figure 2, provides two types of communication paths: i) Cross-lane loopback path (path 3), enabling direct GPU-to-GPU communication within the node, which can be used to construct dynamic size topologies; ii) Dual external paths (path 1&2), connecting to external nodes. All these paths utilize time-division bandwidth allocation, featuring sub-1ms switching latency. With this capability, our OCSTrx allows dynamic reallocation of full GPU bandwidth to an active external path rather than splitting bandwidth across multiple paths. This eliminates redundant link waste—for instance, activating one external path completely disables the other, ensuring efficient bandwidth utilization.
Design 2: Reconfigurable K-Hop Ring topology (§4.2). With OCSTrx  that provides reconfigurable connections at the transceiver, the next challenge is designing the topology. A naive starting point is the the full-mesh topology (Pandey, 2012) which can provide full connectivity among all nodes using OCSTrx . However, full-mesh design requires $O(N^{2})$ links, inducing prohibitive complexity and cost. To reduce costs while maintaining near-ideal fault tolerance and performance, we prune the full-mesh topology into a K-Hop Ring topology based on traffic locality and fault non-locality (Details in §4.2). Combining the reconfigurability of OCSTrx, we propose a reconfigurable K-Hop Ring topology, shown in the middle of Figure 2, which consists of two key parts:
@@
ii) Inter-node fault isolation: dual external paths connect to primary and secondary neighbors (e.g., 2-Hop Ring). When a node fails (e.g., $N_{2}$), its neighbor ($N_{1}$) activates the backup path ($N_{1}$-$N_{3}$) to bypass the fault while maintaining full bandwidth, approaching node-level fault explosion radius. §4.2 generalizes this design to $K>2$.
Design 3: HBD-DCN Orchestration Algorithm (§4.3). Designing an optimal HBD topology is crucial, but end-to-end training performance also depends on the efficient coordination between HBD and DCN. For instance, improper orchestration of TP groups can cause DP traffic to span across ToRs, resulting in DCN congestion. However, existing methods lack the ability to jointly optimize HBD and DCN coordination to alleviate congestion and enhance communication efficiency. To address this, we propose the HBD-DCN Orchestrator, as shown on the right side of Figure 2. The orchestrator takes three inputs: the user-defined job scale and parallelism strategy, the DCN topology and traffic pattern, and the real-time HBD fault pattern. It then generates the TP placement scheme, which maximizes GPU utilization and minimizes cross-ToR communication within the DCN.
@@ 4.1. SiPh-based OCS transceiver (OCSTrx)
The OCSTrx is designed for reconfigurable point-to-multipoint connectivity. It incorporates a compact OCS-based switch with three Rx/Tx paths, utilizing the MZI switch  (Wikipedia, 2024) micro-structure with thermo-optic (TO) effect (Hauffe and Petermann, 2006) phase arms. This OCS-based switch is seamlessly integrated into the Photonic Integrated Chip (PIC) of the transceiver, serving as the MZI switch matrix within the Tx light path, and providing photodetector (PD) modules for each Rx paths.
SiPh-Based OCS. Currently, there are two predominant technological approaches for OCS. Micro Electromechanical systems (MEMS) (Urata et al., 2022a; Yeow et al., 2001) are attractive for commercial adoption due to supporting large port radix, up to a $320\times 320$ matrix (Calient.AI, 2024). Another option is SiPh-based OCS. Its structure is simpler and cheaper to manufacture, the limitation is its radix due to optical losses in multistage light path selector. Given that the locality of traffic and external paths number of OCSTrx is only two, SiPh-based OCS offers greater advantages.
So we choose the design of MZI micro-structure (Wikipedia, 2024) based SiPh-Based OCS. The basic mechanism of MZI switch elements is controlling the phase difference between light paths in two phase arms, then directs the output light to specific ports through interference at the output combiner. TO effect is utilized for phase arm control, for better switching latency compared to MEMS.
OCS Micro-Structure Design. As shown in Figure 3(a), the initial routing decision is made by two MZI switch elements, determining whether to direct the signal through external output 1&2, or the internal loopback path. Subsequently, an internal $N\times N$ MZI switch matrix is incorporated to facilitate the cross-lane loopback mechanism, exemplified by the blue and red paths. Notably, this design can reduces stages count and light attenuation of output 1&2, while ensures consistent light attenuation for them. The design is implemented on the Photonic Integrated Circuit (PIC) chip.
Transceiver Design. In OCSTrx, Tx electrical signal is amplified by linear driver and converted to optical signal by modulators as in Figure 3. One laser is coupled into the PIC as optical source. On the receiving end, multiple photodetectors capture the Rx optical signal from all available paths separately. The output from the activated photodetector is then amplified by a linear transimpedance amplifier (TIA). OCSTrx offers significant benefits, including high compactness, low power consumption, and cost-effective mass production.
@@ 4.2. InfiniteHBD Topology
In this section, we present the InfiniteHBD topology design (Figure 2) integrating OCSTrx that allows all GPUs within datacenter to be connected in a reconfigurable K-Hop Ring topology, while supporting dynamic ring construction and high fault tolerance.
@@
Intra-node Topology. The intra-node topology is designed for dynamic ring construction and compiles with the OCP UBB 2.0 standard ((OAI), 2023). As shown in Figure 4, one node equipped with $R$ GPUs can support $R$ bundles of OCSTrx. Each OCSTrx bundle is connected to a pair of GPUs, with one GPU linking to the upper-half SerDes and the other to the lower-half. For one group of nodes connected as one line, the two GPU paris at each end can interconnect with the OCSTrx internal loopback path, forming a GPU-level ring. As shown in Figure 2, nodes $N_{1}$ and $N_{3}$ are connected in a line, where $OCSTrx_{1}(N_{1})$ and $OCSTrx_{2}(N_{3})$ activate the cross-lane loopback path, creating a ring between the 8 GPUs of $N_{1}$ and $N_{3}$. During ring construction, only two OCSTrx bundles per node are utilized, while the remaining OCSTrx operate in loopback mode. These idle OCSTrx can be replaced with direct connections, such as DAC links, offering a trade-off between cost and reliability. Figure 5(a,b) shows a 4-GPU node with varying numbers of OCSTrx bundles. Note that the topology design in this section utilizes a 4-GPU node as an example, it can be easily scaled for 8-GPU nodes.
@@
Figure 4. OCSTrx connection within nodes. Each block contains multiple OCSTrx as one bundle, .e.g, $8\times 800Gbps$ OCSTrx for a 6.4Tbps GPU.
@@
Inter-node Topology. We construct the inter-node topology by pruning the full-mesh design, based on two key observations: i) Traffic locality: TP Ring-AllReduce in HBD exhibits neighbor communication patterns, eliminating the need for distant connections; ii) Fault non-locality: node-side failures typically occur independently at the node level, meaning consecutive multi-node failures follow an exponentially decaying probability. Each node provides up to $2R$ external paths, allowing us to construct a DC-scale reconfigurable $K$-Hop Ring topology ($K\leq R$) by connecting them to nodes at ±$\pm 1,...,\pm K,K\leq R$. For AllReduce communication, only two out of the $2K$ links are activated once, with the others serving as backup links for fault isolation. For example (Figure 2), if $N_{2}$ fails, $OCSTrx_{2}(N_{1})$ and $OCSTrx_{1}(N_{3})$ can switch to backup links, maintaining connectivity between $N_{1}$ and $N_{3}$ while isolating $N_{2}$’s fault. As $K$ increases, the probability of encountering an unbypassed failure rapidly decreases, which is nearly negligible for $K=3$ (detailed analysis in Appendix §C). Thus, this architecture typically achieves a node-level explosion radius. Moreover, the K-Hop Ring can be broken into the K-Hop line topology, with the trade-off of reduced fault tolerance of $2K$ nodes at two ends.
@@ 4.3. HBD-DCN Orchestration Algorithm
InfiniteHBD is designed to work with arbitrary DCN, including Rail-Optimized (NVIDIA, 2021; Qian et al., 2024) and Fat-Tree (Al-Fares et al., 2008). This section co-optimizes communication performance for both HBD and DCN in InfiniteHBD.
Problem Statement. In InfiniteHBD, GPUs communicate without routing traffic, preventing congestion at any scale. In contrast, DCNs experience inevitable congestion, leading to performance degradation. To mitigate this, we leverage traffic locality to orchestrate nodes, minimizing cross-ToR traffic. Given a job $J$ requiring $N$ nodes from an available pool of $M$ ($M\geq N$), we must select and order $N$ nodes to satisfy two requirements: (1) nodes in the same TP group should communicate via InfiniteHBD, and (2) other parallel traffic should minimize congestion. Ideally, communication remains within the same ToR, confining congestion to switch-to-node links.
@@
Figure 6. Illustration for problem statement of node orchestration.
@@
5. Implementation
OCSTrx: We have successfully built a test board featuring the OCS Controller chip and a pre-release Photonic Integrated Circuit (PIC) module without the MZI switch matrix, as shown in Figure 8. The Controller Chip, measuring $4mm\times 4mm$, is manufactured using a 28nm process, while the PIC, sized at $10.5mm\times 13mm$, uses a 65nm CMOS process. The evaluation board supports 8 pairs of TX/RX SerDes at each end and has been validated for compatibility with various link layer protocols, including PCIe (32Gbps, 64Gbps) and Ethernet (56Gbps, 112Gbps). We assessed the power consumption of the peripheral circuitry using the test board. For an $8\times 112G$ configuration, the power consumption was 8.5 watts. With the addition of 3.2 watts for the MZI switch matrix, the overall consumption totals approximately 12 watts, meeting the QSFP-DD 800Gbps standard(Nowell et al., 2018).
Notably, the full-featured version of the PIC chip has successfully completed tape-out and is currently in the packaging and testing phase. It will be available for evaluation prior to final publication.
Small-scale Cluster: We constructed a small-scale cluster to evaluate the communication performance of the ring topology. Using 32 experimental GPUs equipped with inter-host HBD support (96 lanes on PCIe 4 protocol), we formed a physical ring utilizing fixed optical modules. This mini-cluster was manually reconfigured for both 32-GPU and 16-GPU ring topology. The communication latency and AllReduce performance is evaluated. For small packets, direct GPU-to-GPU links reduced latency by approximately 13% compared to the NVLink switch design. For large packets, the 16-GPU AllReduce utilized 77.11% of the ring bandwidth, with the utilization rate increasing to 77.26% for the 32-GPU configuration, showing minimal degradation with scaling. In comparison, the NVIDIA H100 8-GPU machine achieves an 81.77% utilization rate without SHARP. After deployment of OCSTrx, the size of communication group can be reconfigured within $1ms$, while maintaining maximum throughput.
@@ 6.1. Methodology and Setup
An in-house simulator dedicated for LLM training is used to evaluate InfiniteHBD comprehensively. The simulator supports end-to-end simulations of both model training performance and hardware faults, with the HBD-DCN orchestration algorithm seamlessly integrated into the system.
@@
GPU and network specification. The NVIDIA H100 (NVIDIA, 2024c) (989 TFLOPS, 80GiB) is used for the configuration of GPU in simulation. And the HBD bandwidth of GPU is set as $6.4Tbps$, which is the sum of 8 QSFP-DD OCSTrx. The DCN bandwidth is configured the same as NVIDIA ConnectX-7 ($400Gbps$). Since the simulation primarily focuses on HBD, the DCN is configured as a Fat-Tree topology (Al-Fares et al., 2008). Several HBD architectures are then evaluated, including:
@@
  • NVL-36, NVL-72, NVL-576 (NVIDIA, 2024b): HBDs with 36, 72, or 576 GPUs, GPU are interconnected via NVLink Switches.
@@
GPU count per node. The simulation aligns with both 4-GPU node (.e.g. NVIDIA GB200 NVL-36/72/576 (NVIDIA, 2024b) and TPUv4 (Jouppi et al., 2023)) and 8-GPU node design (NVIDIA H100, AMD MI300X (AMD, 2025), Intel Gaudi3 (Corporation, 2023), and UBB 2.0 standard((OAI), 2023)).
@@
Fault patterns. The fault trace used in the simulation was collected from an 8-GPU node cluster with approximately 10K GPUs over a span of 348 days. On average, the ratio of faulty 8-GPU nodes is $2.66\%$, with the P99 value as $7.92\%$, more details in Appendix §A. In some simulations, fault traces generated based on this trace statistics are also derived.

### v4

@@ 3. Design Overview
Design 1: Silicon Photonics based OCS transceiver (OCSTrx) (§4.1). To enable large-scale deployment, we require a low-cost, low-power transceiver with Optical Circuit Switching (OCS) support. Unlike prior high-radix switches solutions that rely on MEMS-based switching (Urata et al., 2022; Yeow et al., 2001), we leverage advances in Silicon Photonics (SiPh), which offer a simpler structure, lower cost, and reduced power consumption—making them well-suited for commercial transceivers.
Our SiPh-based OCS transceiver (OCSTrx), shown on the left side of Figure 2, provides two types of communication paths: i) Cross-lane loopback path (path 3), enabling direct GPU-to-GPU communication within the node, which can be used to construct dynamic size topologies; ii) Dual external paths (path 1&2), connecting to external nodes. All these paths utilize time-division bandwidth allocation, featuring 60-80 $\mu s$ reconfiguration latency. With this capability, our OCSTrx allows the dynamic reallocation of full GPU bandwidth to a single active external path, rather than splitting it across multiple paths. This eliminates redundant link waste—for instance, activating one external path completely disables the other, ensuring efficient bandwidth utilization.
Design 2: Reconfigurable K-Hop Ring topology (§4.2). With OCSTrx  that provides reconfigurable connections at the transceiver, the next challenge is designing the topology. A naive starting point is the full-mesh topology (Pandey, 2012) which can provide full connectivity among all nodes using OCSTrx . However, full-mesh design requires $O(N^{2})$ links, inducing prohibitive complexity and cost. To reduce costs while maintaining near-ideal fault tolerance and performance, we prune the full-mesh topology into a K-Hop Ring topology based on traffic locality and fault non-locality (Details in §4.2). The topology’s cost and wiring complexity grow linearly with node count, enabling high scalability. Combining the reconfigurability of OCSTrx, we propose a reconfigurable K-Hop Ring topology, shown in the middle of Figure 2, which consists of two key parts:
@@
ii) Inter-node fault isolation: dual external paths connect to primary and secondary neighbors (e.g., 2-Hop Ring). When a node fails (e.g., $N_{2}$), its neighbor ($N_{1}$) activates the backup path ($N_{1}$-$N_{3}$) to bypass the fault while maintaining full bandwidth, thus reducing the fault explosion radius to the node-level. §4.2 generalizes this design to $K>2$.
Design 3: HBD-DCN Orchestration Algorithm (§4.3). Designing an optimal HBD topology is crucial, but end-to-end training performance also depends on the efficient coordination between HBD and DCN. For instance, improper orchestration of TP groups can cause DP traffic to span across ToRs, resulting in DCN congestion. However, existing methods lack the ability to jointly optimize HBD and DCN coordination to alleviate congestion and enhance communication efficiency. To address this, we propose the HBD-DCN Orchestration Algorithm, as shown on the right side of Figure 2. The orchestrator takes three inputs: the user-defined job scale and parallelism strategy, the DCN topology and traffic pattern, and the real-time HBD fault pattern. It then generates the TP placement scheme, which maximizes GPU utilization and minimizes cross-ToR communication within the DCN.
@@ 4.1. SiPh-based OCS transceiver (OCSTrx)
The OCSTrx is designed for reconfigurable point-to-multipoint connectivity. It incorporates a compact OCS-based switch with three Rx/Tx paths, utilizing the MZI switch  (Wikipedia, 2024) micro-structure with thermo-optic (TO) effect (Hauffe and Petermann, 2006) phase arms. This OCS-based switch is seamlessly integrated into the Photonic Integrated Chip (PIC) of the transceiver, serving as the MZI switch matrix within the Tx light path, and providing photodetector (PD) modules for each Rx path.
SiPh-Based OCS. Currently, there are two predominant technological approaches for OCS. Micro Electromechanical systems (MEMS) (Urata et al., 2022; Yeow et al., 2001) are attractive for commercial adoption because they support large port radix, up to a $320\times 320$ matrix (Calient.AI, 2024). Another option is SiPh-based OCS. Its structure is simpler and cheaper to manufacture, the limitation is its radix due to optical losses in the multistage light path selector. Given that the locality of traffic and external paths count of OCSTrx is only two, SiPh-based OCS offers greater advantages.
So we choose the design of an SiPh-based OCS using the MZI micro-structure (Wikipedia, 2024). The basic mechanism of MZI switch elements is controlling the phase difference between light paths in two phase arms, and then directs the output light to specific ports through interference at the output combiner. TO effect is utilized for phase arm control, for lower reconfiguration latency compared to MEMS.
OCS Micro-Structure Design. As shown in Figure 3(a), the initial routing decision is made by two MZI switch elements, determining whether to direct the signal through external output 1&2, or the internal loopback path. Subsequently, an internal $N\times N$ MZI switch matrix is incorporated to facilitate the cross-lane loopback mechanism, exemplified by the blue and red paths. Notably, this design can reduce stages count and light attenuation of output 1&2, while ensuring consistent light attenuation for them. The design is implemented on the Photonic Integrated Circuit (PIC) chip.
Transceiver Design. In OCSTrx, the Tx electrical signal is amplified by linear driver and converted to optical signal by modulators as in Figure 3. One laser is coupled into the PIC as the optical source. On the receiving end, multiple photodetectors capture the Rx optical signal from all available paths separately. The output from the activated photodetector is then amplified by a linear transimpedance amplifier (TIA). OCSTrx offers significant benefits, including high compactness, low power consumption, and cost-effective mass production. Moreover, OCSTrx is integrated into commercial transceivers, and its failures manifest as regular transceiver failures without introducing new failure patterns.
@@ 4.2. InfiniteHBD Topology
In this section, we present the InfiniteHBD topology design (Figure 2) integrating OCSTrx that allows all GPUs within the datacenter to be connected in a reconfigurable K-Hop Ring topology, while supporting dynamic ring construction and high fault tolerance.
@@
Intra-node Topology. The intra-node topology is designed for dynamic ring construction and complies with the OCP UBB 2.0 standard ((OAI), 2023). As shown in Figure 4, one node equipped with $R$ GPUs can support $R$ bundles of OCSTrx. Each OCSTrx bundle is connected to a pair of GPUs, with one GPU linking to the upper-half SerDes and the other to the lower-half. For one group of nodes connected as one line, the two GPU pairs at each end can interconnect with the OCSTrx internal loopback path, forming a GPU-level ring. As shown in Figure 2, nodes $N_{1}$ and $N_{3}$ are connected in a line, where $OCSTrx_{1}(N_{1})$ and $OCSTrx_{2}(N_{3})$ activate the cross-lane loopback path, creating a ring between the 8 GPUs of $N_{1}$ and $N_{3}$. During ring construction, only two OCSTrx bundles per node are utilized, while the remaining OCSTrx operate in loopback mode. These idle OCSTrx can be replaced with direct connections, such as DAC links, offering a trade-off between cost and reliability. Figure 5(a,b) shows a 4-GPU node with varying numbers of OCSTrx bundles. Note that the topology design in this section utilizes a 4-GPU node as an example; it can be easily scaled for 8-GPU nodes.
@@
Figure 4. OCSTrx connection within nodes. Each block contains multiple OCSTrx as one bundle, .e.g., $8\times 800Gbps$ OCSTrx for a 6.4Tbps GPU.
@@
Inter-node Topology. We construct the inter-node topology by pruning the full-mesh design, based on two key observations: i) Traffic locality: TP Ring-AllReduce in HBD exhibits neighbor communication patterns, eliminating the need for distant connections; ii) Fault non-locality: node-side failures typically occur independently at the node level, meaning consecutive multi-node failures follow an exponentially decaying probability. Each node provides up to $2R$ external paths, allowing us to construct a DC-scale reconfigurable $K$-Hop Ring topology ($K\leq R$) by connecting them to nodes at ±$\pm 1,...,\pm K$ ($K\leq R$). Under this design, each node has a degree of $2K$, which is sufficient for traffic locality and tunable via OCSTrx count and intra-node design. For AllReduce communication, only two out of the $2K$ links are simultaneously activated, with the others serving as backup links for fault isolation. For example (Figure 2), if $N_{2}$ fails, $OCSTrx_{2}(N_{1})$ and $OCSTrx_{1}(N_{3})$ can switch to backup links, maintaining connectivity between $N_{1}$ and $N_{3}$ while isolating $N_{2}$’s fault. As $K$ increases, the probability of encountering an unbypassed failure rapidly decreases, which is nearly negligible for $K=3$ (detailed analysis in Appendix §C). Thus, this architecture typically achieves a node-level explosion radius. Moreover, the K-Hop Ring can be broken into the K-Hop line topology, with the trade-off of reduced fault tolerance, affecting the $2K$ nodes at both ends
@@ 4.3. HBD-DCN Orchestration Algorithm
InfiniteHBD is designed to work with arbitrary DCNs, including Rail-Optimized (NVIDIA, 2021; Qian et al., 2024) and Fat-Tree (Al-Fares et al., 2008). This section co-optimizes communication performance for both HBD and DCN in InfiniteHBD.
Problem Statement. In InfiniteHBD, GPUs communicate without routing traffic, preventing congestion at any scale. In contrast, DCNs experience inevitable congestion, leading to performance degradation. To mitigate this, we leverage traffic locality to orchestrate nodes, minimizing cross-ToR traffic. Given a job $J$ requiring $N$ nodes from an available pool of $M$ ($M\geq N$), we must select and order $N$ nodes to satisfy two requirements: (1) nodes in the same TP group should communicate via InfiniteHBD, and (2) other parallel traffic should be arranged to minimize congestion. Ideally, communication remains within the same ToR, confining congestion to switch-to-node links.
@@
Figure 6. Illustration of the node orchestration problem statement.
@@
5. Hardware and Small-Scale Cluster Evaluation
5.1. OCSTrx Hardware Evaluation
Following the design principles outlined in §4.1, we have successfully implemented OCSTrx, a fully integrated module within a QSFP-DD 800Gbps transceiver, as illustrated in Figure 8. As depicted in Figure 9, OCSTrx integrates an OCS Controller Chip and a Photonic Integrated Circuit (PIC) that includes an MZI switch matrix. The Controller Chip, measuring $4mm\times 4mm$, is manufactured using a 28nm process, while the PIC, sized at $10.5mm\times 13mm$, uses a 65nm CMOS process. OCSTrx supports 8 pairs of TX/RX SerDes at each end and has been validated for compatibility with various link layer protocols, including PCIe (32Gbps, 64Gbps) and Ethernet (56Gbps, 112Gbps). OCSTrx has also passed temperature cycling tests, validating its reliability and availability.
(a) Main hardware components.
(b) Packaged module.
Figure 8. OCSTrx integrated in QSFP-DD 800Gbps transceiver.
@@ 5.1. OCSTrx Hardware Evaluation
Reconfiguration latency. Experimental results show that OCSTrx achieves a reconfiguration latency of 60–80 $\mu s$, which is significantly faster than traditional centralized OCS-based switches employing MEMS (Urata et al., 2022), Piezo (Polatis, 2025), or Robotic (Telescent, 2025) mechanisms, whose latencies typically range from milliseconds to minutes. Notably, this measurement excludes software-level delays such as reconnection at the network protocol layer.
Insertion loss. We evaluated the insertion loss of the core module within OCSTrx, which provides the OCS capability, under various ambient temperatures. Figure 10(a) summarizes the statistical results, while Figure 11 shows the distribution. The measured insertion loss ranges from 2.5 dB to 4.0 dB, with an average of 3.3 dB at room temperature (25°C).
Power consumption. Under the $8\times 112G$ configuration, the peripheral circuitry consumes 8.5 $Watts$. We further measured the power consumption of OCSTrx’s core module which provides the OCS capability across different temperatures when activating three paths (two external paths and one internal loopback), as illustrated in Figure 10(b). Across all conditions, the core module consumed less than 3.2 $Watts$, keeping the total power below 12 $Watts$, meeting the QSFP-DD 800Gbps specification (Nowell et al., 2018).
Bit error rate. With the Amplitude Control (AC) driving power set to 620 mW, we measured the bit error rate (BER) of OCSTrx under different optical modulation amplitudes (OMAs) and ambient temperatures, as shown in Figure 12. At -5°C and 25°C, BER was consistently 0. At 50°C and 75°C, BER remained 0 in most cases, with occasional errors only at very low OMAs. These results confirm compliance with the industrial BER threshold.
5.2. Small-Scale Cluster Evaluation
We constructed a small-scale cluster to evaluate the communication performance of the ring topology. Using 32 experimental GPUs equipped with inter-host HBD support (96 lanes on PCIe 4 protocol), we formed a physical ring utilizing fixed optical modules. This mini-cluster was manually reconfigured for both 32-GPU and 16-GPU ring topologies. The communication latency and AllReduce performance are evaluated. For small packets, direct GPU-to-GPU links reduced latency by approximately 13% compared to the NVLink switch design. For large packets, the 16-GPU AllReduce utilized 77.11% of the ring bandwidth, with the utilization rate increasing to 77.26% for the 32-GPU configuration, showing minimal degradation with scaling. In comparison, the NVIDIA H100 8-GPU machine achieves an 81.77% utilization rate without SHARP.
Additionally, a control plane is included to manage the InfiniteHBD. At the device level, the node fabric manager configures individual OCSTrx modules and handles topology switching. At the system level, cluster manager coordinates global control across the cluster. As control-plane mechanisms are not the focus of this work, we omit further details.
(a) Insertion loss.
(b) Power consumption.
Figure 10. Insertion loss and power consumption of the core module in OCSTrx.
(a) 0°C.
(b) 25°C.
(c) 50°C.
(d) 85°C.
Figure 11. Insertion loss distribution of the core module in OCSTrx under different ambient temperatures.
(a) -5°C.
(b) 25°C.
(c) 50°C.
(d) 75°C.
Figure 12. Bit error rate of OCSTrx under varying OMA and ambient temperatures.
@@ 6.1. Methodology and Setup
An in-house simulator dedicated to LLM training is used to evaluate InfiniteHBD comprehensively. The simulator supports end-to-end simulations of both model training performance and hardware faults, with the HBD-DCN orchestration algorithm seamlessly integrated into the system.
@@
GPU and network specification. The NVIDIA H100 (NVIDIA, 2024c) (989 TFLOPS, 80GiB) is used as the configuration of GPU in simulation. The HBD bandwidth of GPU is set to $6.4Tbps$, which is the sum of 8 QSFP-DD OCSTrx. The DCN bandwidth is configured to match the NVIDIA ConnectX-7 ($400Gbps$). Since the simulation primarily focuses on HBD, the DCN is configured as a Fat-Tree topology (Al-Fares et al., 2008). Several HBD architectures are then evaluated, including:
@@
  • NVL-36, NVL-72, NVL-576 (NVIDIA, 2024b): HBDs with 36, 72, or 576 GPUs, GPUs are interconnected via NVLink Switches.
@@
GPU count per node. The simulation aligns with both 4-GPU node (e.g., NVIDIA GB200 NVL-36/72/576 (NVIDIA, 2024b) and TPUv4 (Jouppi et al., 2023)) and 8-GPU node designs (NVIDIA H100, AMD MI300X (AMD, 2025), Intel Gaudi3 (Corporation, 2023), and UBB 2.0 standard ((OAI), 2023)).
@@
Fault patterns. The fault trace used in the simulation was collected from an 8-GPU node cluster with approximately 3K-GPUs over 348 days. On average, the ratio of faulty 8-GPU nodes is $2.33\%$, with the P99 value as $7.22\%$, more details in Appendix §A. In some simulations, fault traces generated based on this trace statistics are also derived.